OpenStudy (anonymous):

Hey can y'all help? What's 8+4-2^2-5=??? First right answer gets a medal!!!!

OpenStudy (anonymous):

12-4-5

OpenStudy (anonymous):

8-5

OpenStudy (anonymous):

depends on how u see it. 8+4 = 12, and -2^2-5 = 4-5 = 9 so 12-9 =3

OpenStudy (anonymous):

=3

OpenStudy (anonymous):

Beat you too it :)

OpenStudy (anonymous):

Okay...um...who do I give the medal to?

OpenStudy (anonymous):

The one with the 58